OpenStudy (anonymous):

Let f(x) = x2 and compute the Riemann sum of f over the interval [3,5], using the following number of subintervals (n). In each case, choose the representative points to be the left endpoints of the subintervals. (Round your answers to two decimal places.) a)Use two subintervals of equal length (n = 2). b)n=5 c) n=10 d) guess the area

OpenStudy (anonymous):

this person asks the same question as you

OpenStudy (anonymous):

it will be of a big help

OpenStudy (anonymous):

ok, so you're gonna approximate the area under the curve using two subintervals...[8,9] and [9, 10]. the width of each of these subintervals is 1.

OpenStudy (anonymous):

since the problem asks for the left endpoint, that will be where your height of the rectangle comes from.

OpenStudy (anonymous):

so with n=2, the approximate area under f(x)=x^2 is the area of those two rectangles.

OpenStudy (anonymous):

What if the intervals are [3,5]

OpenStudy (anonymous):

well you would follow the same steps as given in he link that i gave you, its important that you know so that you can do it for the next time. :)

OpenStudy (anonymous):

I tried working the problem as stated in the link above, I had read that stream before I logged on to ask mine with the interval of [3,5]. I really am lost.

OpenStudy (anonymous):

okay I see well what exactly do you not understand??

OpenStudy (anonymous):

for n=2 I got 1*f(3)=9 and 1*f(4)=16 which 9+16=25 but that is wrong. I honestly don't even know where to start. I started by finding delta x which was 5-3/2 for n=2

OpenStudy (anonymous):

okay I see, im not good at explaining things lol but i will refer jim_thompson5910 he is very good at math and he is online, he explain everything down to a T. you are on the right tract not too far gone but he can help you in the best ways.

OpenStudy (anonymous):

how do i find him

OpenStudy (anonymous):

pluss i mean Im a 10 th grader im not sure what grade your in but am trying my best to explain this lol here i will message him

OpenStudy (anonymous):

okay i just messaged him

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

he is here :)

OpenStudy (anonymous):

dont be sorry you did really good :)

OpenStudy (anonymous):

it me who need to explain things better. lol

jimthompson5910 (jim_thompson5910):

ok one sec, let me read it through

OpenStudy (anonymous):

Well my teacher certainly is not the best and its been a while since I have done this level of Calculus

OpenStudy (anonymous):

hey will im in 10th have never been "taught" calculus i have just leaned a bit on my own lol

OpenStudy (anonymous):

ouch considering I am a Junior in college I REALLY feel pathetic now

OpenStudy (anonymous):

don't lol its no big deal every one needs help :P

jimthompson5910 (jim_thompson5910):

don't feel bad, I didn't learn calc til way late either

OpenStudy (anonymous):

I learned it in high school but don't really remember it all sadly and with my teacher barely able to teach it I am even more lost and she is of no help

OpenStudy (anonymous):

I would really appreciate it if you could help really explain it though Jim

jimthompson5910 (jim_thompson5910):

Ok let's first draw out the parabola y = x^2 on the xy axis  it's supposed to be symmetrical, but oh well lol